# QUART-Online: Latency-Free Multimodal Large Language Model for Quadruped Robot Learning

### v3

@@ III-A Preliminary
$\operatorname{QUART}(\hat{a_{c}}|s,w)=\operatorname{Detokenize}(\hat{a_{d}})p(\hat{a_{d}}|c)\tau(c|s,w).$ (2)
@@ III-B QUART-Online
$\operatorname{QUART-Online}(\hat{A_{c}}|s,w)=\mathcal{D}(\hat{A_{c}}|\hat{A_{q}})p(\hat{A_{q}}|c)\tau(c|s,w),$ (3)
@@ III-C Action Chunk Discretization
Action Encoder. To efficiently encode a greater number of temporal action steps into a more compact format, we use $N_{L}$ layers of 1D convolutions on the temporal axis to compactly encode more action steps, ensuring temporal compression and step coherence. This temporal encoder $\mathcal{E}$ maps action sequences into a high-level latent space, allowing for the extraction of rich semantic and temporal details. Formally, we treat the sequence of $N$ consecutive action commands at time $T$ as a set of input action commands ${A_{c}}=\{a^{i}\}_{i=T}^{T+N}\in\mathbb{R}^{N\times 12}$.
@@
Quantization. Notably, our method utilizes a multi-layer quantization, where the input data is sequentially quantized by each quantizer, and the residuals from each quantization step are explicitly managed. The learnable codebook $\mathcal{B}\in\mathbb{R}^{N_{q}\times K\times D}$ consists of $N_{q}$ layers of codes, where the $i$-th codebook layer are $\mathcal{B}_{i}=\{b^{k}\}_{k=1}^{K}$ including $b^{k}\in\mathbb{R}^{D}$ Here $K$ denotes the number of codes in the codebook, and $D$ represents the dimension of the latent space, corresponding to the number of channels obtained from the encoder. The quantization iterates $N_{q}$ times, culminating in the output of the quantized vector ${A}_{q}\in\mathbb{R}^{C\times N_{q}}$. Specifically, during the $i$-th quantization layer, the input vector of the layer is compared with the corresponding codebook layer $\mathcal{B}_{i}$ to determine the closest match, based on distance calculations. The input vector of each layer is the residual from the previous quantization step, representing the discrepancy between the selected embedding vector and the input vector of that layer. By successively quantifying these residuals, the quantization algorithm refines its approximation of the original high-dimensional data, ensuring encoding accuracy while efficiently managing computational complexity.
@@
Dequantization. The dequantize process is the inverse operation of quantization. Dequantization commences at the apex with layer $\mathcal{B}_{N_{q}}$ and descends to the foundational layer $\mathcal{B}_{1}$. Utilizing indices from $\mathcal{A}_{q}$, it extracts code vectors successively from each layer, aggregating them through a process of accumulation to refine the approximation of the residual incurred at each preceding quantization stage.
@@
In our approach, we leverage a combination of reconstruction loss and adversarial loss for ACD to achieve high-quality quantization and reconstruction of action tokens.
@@ III-D Action Chunk Alignment
Training Phase. Upon obtaining a trained vector quantization representation, the target output for the large model transitions from the original discrete values, which span from 0 to 256, to the compressed action tokens. These tokens, in unison with the codebook and the action decoder, facilitate the inference of the final predicted actions. During this phase, the MLLM undergoes fine-tuning, while the action decoder remains fixed. Given that the final predicted action is a continuous representation, the loss function utilized for this stage follows the cross-entropy loss function employed in the QUART training process.
@@
Inference Phase. Employing a precisely aligned and concise action chunk representation, we effectively minimize the inference time, thereby enhancing the smoothness and accuracy of the generated actions. Additionally, by judiciously selecting an appropriate chunk length, we can achieve latent-free control for large quadruped robot models. This is accomplished when the action chunk length $l_{ac}$ multiplied by the frequency of MLLM inference $f_{m}$ matches the frequency of low-level control $f_{l}$. With this well-calibrated chunk length, we can seamlessly implement latency-free control over large quadruped robot models.
@@ IV-C Comparison with SOTA methods
QUART-Online versus other baseline models.
Compared to other baseline models, QUART-Online demonstrates superior performance across all tasks while maintaining a similar inference rate. This reinforces the effectiveness of our approach, which combines the benefits of large models with the real-time capabilities of lightweight models. Table II illustrates that QUART-Online achieves high success rates, particularly in the "Go avoid" and "Crawl" tasks, outperforming VLA (CLIP) and VLA (VC-1). This combination of high success rates and increased inference frequency enhances the applicability of large models in dynamic robotic tasks.
@@ V Discussion and Limitations
QUART-Online operates in tandem with the existing MLLM system, achieving real-time inference in sync with the underlying controller frequency, significantly boosting the success rate across various tasks. However, there also exist several limitations: 1) The current actions still do not reach the level of joint angles and still require low-level control to obtain joint angle actions. Therefore, exploring how to directly reach the joint angle layer with a large model in the future is a worthwhile area of research.
2) Our rapid-inferring model has yet to be tested in more complex locomotion scenarios. Ensuring that our efficient inference model can adapt to various complex terrains and further tapping into its potential will be the focus of our next phase of work.

### v5

@@ III-A Preliminary
$\operatorname{QUART}(\hat{a_{c}}|s,w)=\operatorname{Detokenize}(\hat{a_{d}})p(\hat{a_{d}}|c)\tau(c|s,w),$ (2)
@@ III-B QUART-Online
$\operatorname{QUART-Online}(\hat{A_{c}}|s,w)=\mathcal{D}(\hat{A_{c}}|\hat{A_{q}})p(\hat{A_{q}}|c)\tau(c|s,w).$ (3)
@@ III-C Action Chunk Discretization
Action Encoder. To efficiently encode a greater number of temporal action steps into a more compact format, we use $N_{L}$ layers of 1D convolutions on the temporal axis to compactly encode more action steps, ensuring temporal compression and step coherence. This temporal encoder $\mathcal{E}$ maps action sequences into a high-level latent space, allowing for the extraction of rich semantic and temporal details. Formally, we treat the sequence of $N$ consecutive action commands at time $T$ as a set of input action commands ${A_{c}}=\{a^{i}\}_{i=T}^{T+N}\in\mathbb{R}^{N\times 12}$:
@@
Quantization. Notably, our method utilizes a multi-layer quantization, where the input data is sequentially quantized by each quantizer, and the residuals from each quantization step are explicitly managed. The learnable codebook $\mathcal{B}\in\mathbb{R}^{N_{q}\times K\times D}$ consists of $N_{q}$ layers of codes, where the $i$-th codebook layer are $\mathcal{B}_{i}=\{b^{k}\}_{k=1}^{K}$ including $b^{k}\in\mathbb{R}^{D}$ Here $K$ denotes the number of codes in the codebook, and $D$ represents the dimension of the latent space, corresponding to the number of channels obtained from the encoder. The quantization iterates $N_{q}$ times, culminating in the output of the quantized vector ${A}_{q}\in\mathbb{R}^{C\times N_{q}}$. Specifically, during the $i$-th quantization layer, the input vector of the layer is compared with the corresponding codebook layer $\mathcal{B}_{i}$ to determine the closest match, based on distance calculations. The input vector of each layer is the residual from the previous quantization step, representing the discrepancy between the selected embedding vector and the input vector of that layer. By successively quantifying these residuals, the quantization algorithm refines its approximation of the original high-dimensional data, ensuring encoding accuracy while efficiently managing computational complexity:
@@
Dequantization. The dequantize process is the inverse operation of quantization. Dequantization commences at the apex with layer $\mathcal{B}_{N_{q}}$ and descends to the foundational layer $\mathcal{B}_{1}$. Utilizing indices from $\mathcal{A}_{q}$, it extracts code vectors successively from each layer, aggregating them through a process of accumulation to refine the approximation of the residual incurred at each preceding quantization stage:
@@
In our approach, we leverage a combination of reconstruction loss and adversarial loss for ACD to achieve high-quality quantization and reconstruction of action tokens:
@@ III-D Action Chunk Alignment
Training Phase. Upon obtaining a trained vector quantization representation, the target output for the large model transitions from the original discrete values, which span from 0 to 256, to the compressed action tokens. These tokens, in unison with the codebook and the action decoder, facilitate the inference of the final predicted actions. During this phase, the MLLM undergoes fine-tuning, while the action decoder remains fixed. Given that the final predicted action is a continuous representation, the loss function utilized for this stage follows the cross-entropy loss function employed in the QUART training process:
@@
Inference Phase. Employing a precisely aligned and concise action chunk representation, we effectively minimize the inference time, thereby enhancing the smoothness and accuracy of the generated actions. Additionally, by judiciously selecting an appropriate chunk length, we can achieve latent-free control for large quadruped robot models. This is accomplished when the action chunk length $l_{ac}$ multiplied by the frequency of MLLM inference $f_{m}$ matches the frequency of low-level control $f_{l}$. With this well-calibrated chunk length, we can seamlessly implement latency-free control over large quadruped robot models:
@@ IV-C Comparison with SOTA methods
QUART-Online versus other baseline models. Compared to other baseline models, QUART-Online demonstrates superior performance across all tasks while maintaining a similar inference rate. This reinforces the effectiveness of our approach, which combines the benefits of large models with the real-time capabilities of lightweight models. Table II illustrates that QUART-Online achieves high success rates, particularly in the "Go avoid" and "Crawl" tasks, outperforming VLA (CLIP) and VLA (VC-1). This combination of high success rates and increased inference frequency enhances the applicability of large models in dynamic robotic tasks.
@@ V Discussion and Limitations
QUART-Online operates in tandem with the existing MLLM system, achieving real-time inference in sync with the underlying controller frequency, significantly boosting the success rate across various tasks. However, there also exist several limitations: 1) The current actions still do not reach the level of joint angles and still require low-level control to obtain joint angle actions. Therefore, exploring how to directly reach the joint angle layer with a large model in the future is a worthwhile area of research. 2) Our rapid-inferring model has yet to be tested in more complex locomotion scenarios. Ensuring that our efficient inference model can adapt to various complex terrains and further tapping into its potential will be the focus of our next phase of work.
@@ VI Conclusion
Acknowledgements
This work was supported by the National Science and Technology Innovation 2030 - Major Project (Grant No. 2022ZD0208800), and NSFC General Program (Grant No. 62176215).